x/3+ x = 8
and
4/x +3 = 1/x


I must show my work and i don't know how ???? please help me !!!!!

Answer by edjones(8007)

x/3 + x = 8 I assume that this is the problem.
x+3x=24 multiply each side by the LCM to eliminate fractions.
4x=24
x=6
.
4/x + 3 = 1/x
4+3x=1 multiply each side by the LCM to eliminate fractions.
3x=-3
x=-1
.
Ed